Two students at Sky High School will be chosen for a TV interview. If there are 900 students at school, how many different selections are possible?
Answer by Edwin McCravy(20059)


Two students at Sky High School will be chosen for a TV interview. If there are
900 students at school, how many different selections are possible?
                 
C(900,2) = (900·899)/(2·1) = 404550

Twenty people report for jury duty. How many different 12-person juries can be
chosen?

C(20,12) = (20·19·18·17·16·15·14·13·12·11·10·9)/(12·11·10·9·8·7·6·5·4·3·2·1)
= 125970

How many different committees of three members can be formed from a group of
eight people?
	
C(8,3) = (8·7·6)/(3·2·1) = 56
	
Edwin